OpenStudy (anonymous):

Write an equation of a line containing the given point and the parallel to the given line. express your answer in the form y==mx+b. (6,9); x+5y=8

OpenStudy (anonymous):

When two lines are parallel it means they have the same slope. What is the slope of x+5y=8?

OpenStudy (anonymous):

Try to get y on side.

OpenStudy (anonymous):

\[y = -x\frac{1}{5} + \frac{8}{5}\]

OpenStudy (anonymous):

So the slope of this line and our line is -1/5

OpenStudy (anonymous):

So we already know the m=-1/5

OpenStudy (anonymous):

We can find the b,"usually known as the y intercept" by pluging in (6,9) into y = (-1/5)x + b. =]

OpenStudy (anonymous):

-1/5x+?/5 I can figure out how to sole for the ?

OpenStudy (anonymous):

Forget about the first line. It already gave us what we wanted: The slope. Our line is running parallel to it so its slope is the same. If you can imagine two lines apart running in the same direction and will never cross. So OUR line is like this: y = (-1/5)x + b We must find b. We can do that because we know a point on it. We put in x=6 and y=9. Then get b on one side.

OpenStudy (anonymous):

i got 5.8. is that right

OpenStudy (anonymous):

9b=-5.8

OpenStudy (anonymous):

Show me your work. does (-1/5) * 6 = (-6/5) = -5.8?

OpenStudy (anonymous):

(9) = -(6/5) + b

OpenStudy (anonymous):

9 + (6/5) = b

OpenStudy (anonymous):

Yummy fractions.

OpenStudy (anonymous):

b = 51/5 or 10.2. You probably got the same thing.

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

Just remember TWO parallel lines have the same slope, the same m. Good luck! =]